OpenStudy (anonymous):

A solution may be a homogeneous or heterogeneous mixture of solute in solvent. True or False

OpenStudy (anonymous):

F

OpenStudy (anonymous):

Solution - homogeneous mixture of 2 or more components

OpenStudy (somy):

i think False also

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which solution is most concentrated: 200 ppm 20 ppt 2 % All of the above have same concentration

OpenStudy (anonymous):

200ppm I believe

OpenStudy (somy):

ugh damn conversion

OpenStudy (anonymous):

– ppt = (g solute / g solution) * 10^3 ppt – ppm = (g solute / g solution) * 10^6 ppt

OpenStudy (anonymous):

The most concentration is the one that has the most amount right

OpenStudy (somy):

i guess so

OpenStudy (anonymous):

Tyndall effect is passage of light through a solution without scattering the ability of a suspension to scatter the light (hazing effect) Creation of white color in solution settlement of solid particle from a solution

OpenStudy (anonymous):

the ability of a suspension to scatter the light (hazing effect)

OpenStudy (somy):

is light scattering by particles in a colloid or particles in a fine suspension

OpenStudy (somy):

so yeah u r right i think

OpenStudy (anonymous):

What is the molarity of a 5L solution that contains 2.5 moles of HNO3? 5 Molar 2.5 Molar 0.5 Molar 12.5 Molar

OpenStudy (somy):

molarity= mole/volume

OpenStudy (anonymous):

so is 0.5

OpenStudy (somy):

yes

OpenStudy (anonymous):

You have 50 mL of a 0.1M aq. soln. of HCl and have to dilute it to a final volume of 1L. What is the molarity of final solution? 0.01M 0.1 M 50 M 0.005 M

OpenStudy (somy):

C1V1= C2V2 C1 = 0.1 V1 = 50mL /1000= 0.05L C2 = X V2 = 1 L

OpenStudy (anonymous):

.005

OpenStudy (somy):

yes

OpenStudy (anonymous):

18 mins left, 6 questions left

OpenStudy (anonymous):

Solution properties that depend on the concentration of particles rather than the chemical identity and concentration of solute molecules are called ‘Colligative Properties’ True False

OpenStudy (anonymous):

Colligative properties - properties of solutions that depend on the concentration of the solute particles, rather than the identity of the solute

OpenStudy (somy):

as u can see its true

OpenStudy (anonymous):

Which of the following will do the maximum elevation of boiling point or depression in freezing point? 1 mol of NaCl 1 mol of Na3PO4 1 mol of Na2SO4 1 mol of sugar

OpenStudy (anonymous):

NaCl

OpenStudy (somy):

idk this one T_T

OpenStudy (anonymous):

What is the osmotic pressure of 0.5 molar NaCl solution at 25 oC? (R = 0.0821 L.atm/K.mol); pi = MRT (T in Kelvin) 2.4 atm 2.05 atm 24.5 atm Need more data to calculate

OpenStudy (somy):

what is this --> pi = MRT

OpenStudy (anonymous):

Yeah I get 11.0835. Im doing something wrong

OpenStudy (somy):

o.5*0.0821* (25+273)

OpenStudy (anonymous):

12.2329

OpenStudy (somy):

tcheh weird T_T

OpenStudy (anonymous):

what is 25C in kelvins

OpenStudy (anonymous):

isn't 273 + 298?

OpenStudy (somy):

no its 25+ 273

OpenStudy (anonymous):

Colloidal suspensions contain solute particles of 1 nm -200 nm size that are not uniformly distributed. True False

OpenStudy (anonymous):

T

OpenStudy (somy):

pi = iMRT !! idk whats i tho

OpenStudy (anonymous):

How many grams of glucose are needed to make its 15% (m/V) solution in 750 mL of water? 15 g glucose 750 g glucose 150 g glucose 112.5 g glucose

OpenStudy (somy):

and yeah i think the previous one is true

OpenStudy (anonymous):

5mins left

OpenStudy (somy):

ugh 15/100= 0.15 m/V i think 0.15= m/V m= 0.15*V

OpenStudy (anonymous):

A saturated solution is a dynamic equilibrium between the process of dissolution and precipitation True False

OpenStudy (somy):

so im getting 112.5 g

OpenStudy (anonymous):

False

OpenStudy (anonymous):

For proper cell functions to happen, cells need to be maintained in Hypertonic solutions Isotonic Solutions Hypotonic solutions any of the above

OpenStudy (anonymous):

Isotonic Solutions

OpenStudy (somy):

yes

OpenStudy (anonymous):

Done

OpenStudy (anonymous):

I keep messing up 3 always

OpenStudy (somy):

ugh >.>

OpenStudy (somy):

3 mistakes?

OpenStudy (anonymous):

yes

OpenStudy (somy):

what are they?

OpenStudy (somy):

in which question

OpenStudy (anonymous):

it doesn't show until wednesday

OpenStudy (somy):

ughhh

OpenStudy (somy):

3 wrong out of?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

wait. That is wrong

OpenStudy (somy):

20% is wrong T_T

OpenStudy (somy):

ai?

OpenStudy (anonymous):

If there are only 12 questions, why it says out of 15 unless some are 2 points

OpenStudy (somy):

o.o

OpenStudy (anonymous):

yes

OpenStudy (somy):

whatever still 20% is lost hhh

OpenStudy (anonymous):

I know. I can't make above 90

OpenStudy (somy):

T_T im sorry

OpenStudy (anonymous):

thats ok.

OpenStudy (anonymous):

I'll go to eat

OpenStudy (somy):

aye..